OpenStudy (anonymous):

Find the length of segment BD. B=-5 D=3

Directrix (directrix):

Is there a diagram?

Directrix (directrix):

Are B and D points on a number line?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

Directrix (directrix):

Looking at the uploaded file below, how far is it from -5 to zero?

Directrix (directrix):

You count the units. We're taking turns solving the problem. And, it is your turn.

OpenStudy (anonymous):

5

Directrix (directrix):

The rules of this site do not allow me to work the problem for you. I'm supposed to work the problem with you. So, please help with that. It's okay if you post a wrong answer. It's okay if you post a correct answer.

Directrix (directrix):

Good. My turn. How far is it from zero to 3. Answer: 3 units. BD = 5 + 3 = 8 units. Now, take a look at the Ruler Postulate.

OpenStudy (anonymous):

substitue it in

Directrix (directrix):

The distance between two points on a number line can be found by taking the absolute value of the difference of their coordinates. We could have done this problem by using this portion of the Ruler Postulate. abs [ (-5) - (3) ] = ? Your Turn.

Directrix (directrix):

You could calculate this first and then take the absolute value: [ (-5) - (3) ] = ? Your Turn

OpenStudy (anonymous):

8

Directrix (directrix):

(-5) - (3) = -5 - 3 = -8. So, the absolute value of (-8) is +8 or 8 which is the distance from B to D. BD = 8 as we counted. Hey, a lot of people do -5 + 3 = -2 and then call 2 the length of BD. Common error.

OpenStudy (anonymous):

so is it 8

Directrix (directrix):

Yes.